Project

General

Profile

RunPlan20150619 » History » Version 3

Drielsma, Fran├žois, 19 June 2015 16:57

1 1 Blackmore, Victoria
h1. June 19th, 2015 (Evening Run)
2
3
*Aim:* Tracker alignment data
4
5
h3. Daily schedule
6
7
| June  | 19th |   
8
|Time |  10pm - 6am |
9
| MOM       | V. Blackmore   |  
10
| Lead shifter | C. Heidt | 
11
| Shifter 2 |  P. Francini  | 
12
| BLOC      | P. Hanlet   | 
13
| SOC       | C. Heidt   | 
14
| Experimenter | E. Overton, M. Uchida  | 
15
| Physics Shifter | C. Rogers  |
16
|  | | | 
17
| Beam?  | Yes  | 
18
| Additional Personnel | | |  
19
20
21
h2. Shift Monitoring
22
23
Shifters should keep a *careful eye on*:
24
* *Water chiller temperature* outside the control room (bottom of the red LEDs).  If this reaches 25C, immediately power down the magnets and inform the MOM.
25
* *Magnet water temperatures.*
26
** Log any abnormal changes in the eLog
27
** If yellow temperature alarm sounds, immediately power down the magnets and inform the MOM.
28
* *Magnet water flows.*
29
** Log any abnormal changes in the eLog
30
** If yellow alarm sounds, immediately power down the magnets and inform the MOM.
31 2 Blackmore, Victoria
* *EMR stopping the DAQ arming*
32 3 Drielsma, Fran├žois
** Power cycle the VME crate by following the instructions "here":http://micewww.pp.rl.ac.uk/projects/operations/wiki/EMR_Detector
33 1 Blackmore, Victoria
34
h2. Shift plan:
35
36
* 8:45pm
37
** Shifters + BLOC arrive
38
** Hall and beam line to be closed and prepared for beam
39
* 10pm
40
** Check magnet chilled water temperatures in archiver
41
** Take 500 spills of Configuration 1 and calculate ratio of TOF1 to TOF2 triggers (i.e. attempt to quantify transmission)
42
** Take 500 spills of Configuration 2 and calculate ratio of TOF1 to TOF2 triggers (i.e. attempt to quantify transmission)
43
** Continue data taking of beam setting with largest transmission, minimum 50k accepted triggers at *TOF2*
44
** If transmission of beams is similar and 50k triggers at TOF2 is easily achieved within shift, take 50k triggers at TOF2 of the alternative configuration.  Otherwise, prefer higher momentum (configuration 1)
45
46
h2. Data Requested
47
48
* (3, 300) muon beam 
49
* Trigger from TOF1
50
* 50k triggers at *TOF2*
51
52
53
54
h2. Beamline Configurations
55
56
* We expect to run with 2V beamloss and the beam bump active.
57
* DS is off
58
* Beam settings via Run Control
59
60
h3. Configuration 1
61
62
Setting for 300 MeV/c however upstream is set to the max available current (a bit lower than typically coming from Marco's table):
63
64
| Number of pulses| Beamloss |  Particle Species| p at Tgt | p@D1| p@D2| Proton Absorber | Q1 | Q2 | Q3 | D1| DS | D2| Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s)|RunControl tag|
65
| #  | V  |  | MeV/c| MeV/c| MeV/c| mm| A| A| A| A | A | A | A | A|A|A|A|A| DAQ run #||
66
|  |  | Pi+/Mu+ | 512 | 508.75 | 333 | 82  | 109.21 | 199.75 | 122.30 | 437.83| -| 124.71| 221.83 |297.48 | 197.36| 202.43| 306.46 | 261.93|  |300+No_DS|
67
68
h3. Configuration 2
69
70
Setting for 270 MeV/however upstream is set to the max available current, which than corresponds to about 270 MeV/c into the Cooling Channel:
71
72
| Number of pulses| Beamloss |  Particle Species| p at Tgt | p@D1| p@D2| Proton Absorber | Q1 | Q2 | Q3 | D1| DS | D2| Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s)|RunControl tag|
73
| #  | V  |  | MeV/c| MeV/c| MeV/c| mm| A| A| A| A | A | A | A | A|A|A|A|A| DAQ run #||
74
|  |  | Pi+/Mu+ | 512 | 508.75 | 299.26 | 82  | 109.21 | 199.75 | 122.30 | 437.83| -| 112.49| 199.23 |267.19 | 177.23| 180.04| 272.52 | 232.88|  |270+No_DS|
75
76
77
h2. Phone details for 6pm daily meeting
78
79
Phone access :  ReadyTalk number 
80
81
US 866-740-1260
82
83
UK 0800-496-0576
84
85
Swiss 0800-70-53-51
86
87
Italy 800182592
88
89
Access Code: 2155402